P3-45. Compute the DuPont Disaggregation of ROE Refer to the balance sheets and income statement below for Facebook Inc.FACEBOOK INC. Consolidated Statement of Income For Year Ended December 31, $ millions 2018 $55,838 9,355 10,273 7,846 3,451 RFACEBOOK INC. Consolidated Balance Sheet At December 31, $ millions 2018 2017 $10,019 31,095 7,587 1,779 $ 8,079 33,632 5,832Required a. Compute return on equity (ROE). b. Apply the DuPont disaggregation into return on assets (ROA) and financial leve
Answer #1

Solution:

a).

ROE=\frac{Net Income}{Shareholder Funds or Equity}

=\frac{\$22112 }{(\$84127+\$74347)/2}=27.9\%

ROA=\frac{Net Income }{Avg.TotalAssets}

=\frac{22112}{(97334+84524)/2}=24.3\%

Financial Leverage (FL)=\frac{Avg Assets}{AvgShareholder'sEquity}=\frac{90929}{79237}=1.148

AssetsTurnover=\frac{Sales}{AvgAssets}=\frac{\$55838}{90929}=0.614

b).

Dupont disaggregation into ROA and inancial Leverage=ROE=ROA*FL

=24.3%*1.148=27.90%

c).

Profitability and Productivity components of ROA

Profit Margin Shoes thee Profitability of the Company.

Profit Margin=\frac{NetIncome}{Sales}=\frac{\$22112}{55838}=39.6\%

AssetsTurnover shows productivity of the Company.

Asset turnover=\frac{Sales}{AvgAssets}=\frac{\$55838}{90929}=0.614=39.6\%\times0.614=24.3\%

d).

Full DUpont DIsaggregation

=ROE=PM*AT*FL

=39.6%*0.614*1.148

27.9%